A sample of 0.892 g of potassium chloride (KCl) is dissolved in 54.6 g of water. What is the percent by mass of KCl in this solution?
Calculate the percent by volume of 20 mL of alcohol diluted in water to produce 250mL solution.
1
Expert's answer
2020-10-21T12:48:36-0400

1) "%=frac\{};{}""Cm (KCl)=\frac{m(solute)}{m(solute)+m(solvent)}⋅100%"

"%=frac\{};{}""Cm (KCl)=\frac{0.892}{0.892+54.6}⋅100 = 1.6 percent"


2) "Cv =\frac{v(solute)}{v(solute)+v(solvent)}⋅100%"

"Cv (alcohol)=\frac{20}{20+250}⋅100=7.4 percent"